The interest rate of a Japanese one‐year bond is 0.5%, in Australia it is 3.5%. The spot market exchange rate between the currencies of these two countries is 80 Yen/A$. What swap rate can you expect for 6 months and what will be the forward rate for this time horizon?
Answer #1

We know that

Forward Rate=Spot Rate X (1+Interest rate of overseas country)/ (1+interest rate of domestic country)

Also,Interest rate given is for one year and hence we need to divide it by 2 to get 6 months interest rate.

So here

Forward rate=80* (1+0.005/2)/ (1+0.035/2)

=> Forward rate=80* (1.0025/1.0175)=78.82064

Swap rate=80-78.82064=1.179

So Forward rate=78.82064